What's the probability that a sum of 7 or 11 appears at least twice, if the a pair of fair dice are tossed 3 times.
Answer by edjones(7311)

The probability of 11 on one roll is 2/36
The probability of 7 on one roll is 6/36
The probability of 7 or 11 on one roll is 8/36=2/9
(a+b)^3=a^3+3a^2b+3ab^2+b^3
Let a=2/9 and b=7/9
We are only interested in the 1st 2 terms:
(2/9)^3+3(2/9)^2(7/9)=.1262 answer.
.
Ed